if f(x)=6x+7
f(r-2)
i dont know how to plug in the numbers??

Answer by checkley77(12844)

if f(x)=6x+7
f(r-2)
Just substitute (r-2) for x and solve for r.
f(r-2)=6(r-2)+7
f(r-2)=6r-12+7
f(r-2)=6r-5 ans.